Write an equation in slope intercept form for the line that passes through the given point and is parallel to the graph of (-2,4), y= -3x + 10
Answer by algebrahouse.com(1659)

Slope-intercept form is y = mx + b
m is the slope
b is the y-intercept

For y = -3x + 10
slope = -3

Parallel lines have equal slopes, therefore the slope of the new line will be -3.
y - y1 = m(x - x1) {point-slope form}
(x1,y1) is the point and m is the slope
y - 4 = -3[x - (-2)] {substituted point and slope into point-slope form}
y - 4 = -3(x + 2) {simplified double negative signs}
y - 4 = -3x - 6 {used distributive property}
y = -3x - 2 {added 4 to each side}

For more help from me, visit:
www.algebrahouse.com